### v1

# Personality Pairing Improves Human-AI Collaboration

Abstract
Here we ask how AI agent “personalities” interact with human personalities, and other traits, to shape human-AI collaboration, productivity and performance. To estimate these relationships, we conducted a large-scale preregistered randomized experiment that paired 1,258 participants with AI agents that were prompted to exhibit varying levels of the Big Five personality traits. These human-AI teams produced 7,266 display ads for a real think tank, and the quality of these ads was evaluated by 1,995 independent human raters as well as in a field experiment conducted on X, which generated nearly 5 million impressions. We found, first, that personality pairing impacted teamwork quality. For example, neurotic AI improved teamwork for agreeable humans but impaired it for conscientious humans. Second, we found productivity effects of personality pairing and a “productivity-performance trade-off" in which certain pairings (e.g., agreeable human with neurotic AI) produced fewer ads but of higher quality. Third, personality pairing influenced ad quality and performance. For example, quality improved when open humans were paired with conscientious AI and when conscientious humans were paired with disagreeable AI. Some of these pairing effects were “jagged” in that they varied across text and visual tasks. For example open humans produced higher quality images but lower quality text when paired with agreeable AI. Pairing effects were also present in other human traits, like country of origin. For example, extroverted AI improved quality for Latin American workers, but degraded quality for East Asian workers. These findings demonstrate that human-AI personality alignment significantly improves collaboration, productivity, and performance and lay a foundation for future research on improving human-AI collaboration through AI personalization.
Introduction
Artificial intelligence (AI) tools have gained significant attention for their ability to increase productivity and performance across a wide range of tasks. For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (noy2023experimental). AI-enhanced resumes increased job hiring rates by 8% (wiles2023writing). In customer support settings, AI assistance increased productivity by an average of 14% (brynjolfsson2023genai). But these benefits also exhibited considerable heterogeneity, varying across tasks (dellacqua2023navigating, ju2025collaborating), worker seniority (wang2023friend), gender (dellacqua2023navigating, humlum2024adoption), and skill levels (noy2023experimental, brynjolfsson2023genai, otis2024uneven). In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (chen2024impact, yu2024heterogeneity, agarwal2023combining). In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (yang2025my, vaccaro2024combinations, sergeyuk2025human). Unfortunately, experiments documenting these productivity and performance gains have generally provided participants with identical AI tools, leaving open the possibility that at least some of the observed variation in performance stems from a mismatch between AI and individual human characteristics. Performance gains may therefore increase and simultaneously vary less by age, gender, or skill level, if AI is personalized for each human collaborator.
Such personalization may now be possible as foundation model companies have recently begun developing, refining and deploying AI personalities to improve human-AI interaction. For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities (openai2025customizing, openai2025sycophancy). Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences (anthropic2024character, sharma2025sycophancy), or other “evil” traits, and introduced novel training methods to mitigate such personality shifts in their models. These advances enable AI agents to be tailored to users, potentially improving human-AI collaboration and ushering in a new era of AI personalization.
Despite the development of AI personalities, however, no prior research has rigorously and experimentally examined their causal effects on human-AI collaboration. Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork, productivity, and performance as the lack experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, AI psychoses.
In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (zhang2023teammate, hemmer2023, woolley2023collective). Humans exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (kichuk1997, peeters2006, woolley2010ci). Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (barrick1998gma, bell2007meta). As AI has evolved from a passive tool to an active collaborator (makarius2020rising, fugener2022cognitive, anthony2023, collins2024building, ju2025collaborating), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time. Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (li2025acceptance, tully2025lower, fang2025psychosocial).
Unfortunately, although human and AI traits shape human-AI teamwork, their interaction remains understudied (krakowski2025human). With AI agents increasingly operating as social actors in collaborative workflows (anthony2023, gelbrich2025), it is crucial to understand whether the principles of personality alignment apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts. To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork, productivity, and performance.
Our study leverages the unique capabilities of AI agents to vary their individual personalities (jiang2023evaluating, chen2025persona), allowing us to systematically investigate how human-AI personality alignment influences the productivity and performance of human-AI teams. Systematic prompt variation can effectively shape AI agent behavior (schulhoff2024prompt), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (brown2020language). Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (wei2023chainofthought) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (chen2025persona). These methods enable us to intentionally design AI agents with particular social and behavioral characteristics, which can be tailored to complement human collaborators. Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (jakesh2023opinionated).
Investigating human-AI collaboration with agentic tools is challenging in practice. These tools must be seamlessly embedded in workspaces, have full contextual awareness, and be capable of taking actions on behalf of users. Developing such sophisticated tools is a complex and costly endeavor, and rigorous research in this area requires multiple treatment arms to test interactions with individual covariates, necessitating large sample sizes to ensure statistical power. Downstream performance measures are also crucial to evaluating the effectiveness of real collaborations in the field, not just in the lab. Our study addresses all of these requirements, providing a comprehensive framework for investigating the real-world economic impact of AI personalization.
Experiment Design
To investigate how personality alignment between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S. population using Pairit, an experimental platform we developed for extensible human-AI collaboration (ju2025collaborating). After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (jiang2023evaluating). These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank within a 40-minute session, using real-time chat and synchronized text and image editing tools. Teams were provided with image assets by the think tank and could separately call Dall-E 3 to generate new images for the campaigns. As described in the Materials and Methods, the AI agents were fully multi-modal, had access to the Internet, could manipulate the images, the image carousel and the text of the ads, text with their human collaborators and could also call Dall-E 3 to generate new images. This setup allowed us to assess the causal effects of personality alignment on collaboration outcomes. The randomization of AI personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration, productivity and performance of human-AI teams.
Figure 1: The experimental design for investigating human-AI collaboration. (A) Participants are randomized into collaborating with either a human partner or an AI agent. (B) AI agents are randomly assigned prompts to induce low or high levels of the Big Five personality traits. (C) Pairs collaborate in a real-time collaborative workspace to create display ads.
Importantly, we approach personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (chhikara2025mem0, sharma2025sycophancy). Instead, we randomize AI personalities to learn the causal effects of various personality pairings on human-AI team performance. This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (jung2013backchanneling, jung2018theories).
To assess the quality of the display ads produced by the human-AI teams, we conducted a separate evaluation study using a second independent sample of 1,198 new participants recruited from Prolific. These raters were based in the US and selected to ensure representation across gender, age, and ethnicity. Each rater evaluated a random sample of 40 ads, with the sampling design guaranteeing that every ad received at least three independent ratings. Raters assessed each ad on three dimensions using 7-point Likert scales: text quality (“The text is present, clear, relevant, and engaging”), image quality (“The image is visually appealing”), and likelihood of clicking the ad (“I am likely to click on this ad”). The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Supporting Information [SI] Section C).
To evaluate the real-world performance of the display ads, we then conducted a pre-registered field experiment on the X platform. We sampled 2,000 ads from the total 7,266 ads using stratified sampling based on predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution. These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (braun2024abtesting, braun2024adplatforms). Key performance metrics included the click-through rate (CTR) and the cost-per-click (CPC) measured on X, as well as the view-through rate (VTR; fraction of document viewed) and the view-through duration (VTD; seconds) tracked via unique DocSend links for each ad (details of the field experiment are provided in the Materials and Methods).
Results
Our findings revealed that personality pairings significantly affected collaborative outcomes including teamwork quality, productivity, ad quality and ad performance. For example, humans high in conscientiousness created ad text of greater quality when paired with conscientious AI agents, while extroverted humans paired with conscientious AI produced lower quality text, lower quality images, and ads with lower click likelihood. Some pairings demonstrated uneven “jagged” performance patterns, with effects varying sharply between the textual and visual components of the task. We further observed effects on productivity, with agreeable humans, for example, producing more ads when paired with open AI. Additionally, certain combinations yielded a performance-productivity trade-off, such as agreeable humans with neurotic AI generating fewer ads, but ads of higher quality. Overall, these results underscore the substantial role of personality alignment in human-AI teamwork and lay a foundation for future research on AI personalization to improve human-AI productivity and collaborative performance. In particular, we found specific effects on teamwork, productivity and output quality.
Teamwork
First, we found personality pairings significantly influenced teamwork quality. To evaluate these effects, participants completed a 35-item teamwork quality survey after the task, which assessed six facets of collaboration: communication, coordination, balance of member contributions, mutual support, effort, and cohesion (hoegl2001). We conducted a regression analysis on data produced by our experiments to identify the average treatment effects of AI personalities as well as the interaction effects of human and AI personality traits in combination and found statistically significant effects for human openness, conscientiousness, and agreeableness, as well as for AI neuroticism (Figure 2, see SI Section H).
Figure 2: Regression coefficients for human-AI personality interaction terms for teamwork quality (p<<0.05) (hoegl2001). Error bars show 95% confidence intervals. See SI Section H for the full regression tables.
In examining interaction effects, we found that for open humans, conscientious AI lowered the balance of member contributions, mutual support, effort, and cohesion. For conscientious humans, open AI improved balance and cohesion, and conscientious AI improved effort and cohesion. However, for conscientious humans, neurotic AI lowered communication, coordination, balance, and mutual support. Finally, for agreeable humans, neurotic AI improved communication, coordination, mutual support, and effort. However, these positive results reversed when agreeable humans were paired with open AI (for cohesion) and with conscientious AI (for coordination, balance, and cohesion). These effects revealed that personality pairings shape perceived teamwork dynamics in human-AI collaboration.
Ad quality
In human evaluations of ad quality, we found that personality pairing between humans and AI agents significantly influenced ad text quality (Figure 3) beyond the effects of AI personality alone (see SI Section H). The average treatment effects demonstrated that conscientious AI reduced, and extroverted AI improved, text quality on average. The interaction effects then demonstrated statistically and economically significant effects of human-AI personality pairing on ad quality. For both open and conscientious humans, conscientious AI increased text quality, but agreeable AI decreased text quality. For extroverted humans, extroverted AI increased text quality, but conscientious AI decreased text quality. For agreeable humans, neurotic AI increased text quality, but extroverted AI decreased text quality. These results indicate that aligning AI personality traits with human traits impacts the effectiveness of collaborative ad text production.
Figure 3: Regression coefficients for human-AI personality interaction terms on ad text quality, image quality, and click likelihood (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.
For image quality, we observed a different set of significant interaction effects based on human-AI personality pairing. For open humans, agreeable AI increased image quality while, for conscientious humans, open AI increased image quality. For extroverts and agreeable participants, being paired with conscientious and open AI, respectively, decreased image quality.
For click likelihood ratings, which reflect the perceived engagement potential of the ads, we observed no significant effects for humans with high openness or agreeableness. For conscientious humans, agreeable AI decreased click likelihood, while for extroverted and neurotic humans, conscientious AI decreased click likelihood.
Across the three measures of ad quality (text quality, image quality and perceived click likelihood), personality pairings between humans and AI agents consistently influenced outcomes. For extroverts, conscientious AI decreased quality across all measures. For conscientious humans, agreeable AI decreased text quality and click likelihood. Some pairings had significant and opposite effects, indicating a “jagged edge” of human-AI competence across tasks and personality pairings (dellacqua2023navigating). For example, for open humans, agreeable AI improved image quality but worsened text quality. Collectively, these patterns suggest that strategic personality pairings can optimize human-AI collaboration outcomes, with some generalizable and some task-specific personality pairings.
In addition to participant personality traits, we found that participant demographics, including previous AI experience, employment status, and country of origin, moderated the effects of AI personality traits on ad quality. For individuals with previous AI experience, extroverted AI hurt overall performance, while neurotic AI improved performance (see SI Sections G).
Although all participants were based in and representative of the demographic composition of the U.S., their diverse countries of birth allowed us to explore possible cultural and regional differences in collaborative outcomes (Figure 4, see SI Section H). Participants from Western regions, the Middle East & South Asia produced significantly lower quality images when paired with agreeable AI, while participants from East Asia produced images of significantly lower quality when paired with extroverted and agreeable AI. Participants from Latin America produced significantly higher quality ad text when paired with extroverted AI and significantly lower quality ad text when paired with conscientious, agreeable, and neurotic AI. For participants from all regions except East Asia, agreeable AI significantly lowered click likelihood. For those from Latin America, extroverted AI improved click likelihood, while conscientious AI lowered click likelihood. For those from the Middle East & South Asia, agreeable AI lowered click likelihood.
Interestingly, we observed that agreeable AI consistently undermined performance across all measures, reducing image quality (for Western, East Asian, Middle Eastern & South Asian workers), text quality (for Latin American workers), and click likelihood (for Western, Latin American, Middle Eastern & South Asian workers). Taken together, these patterns suggest individual characteristics like culture mediate how AI personality traits, especially agreeable or sycophantic AI (sharma2025sycophancy), affect collaboration and suggest the potential for future research investigating the strategic pairing of AI personalities with various human traits.
Figure 4: Regression coefficients for interaction terms of a participants’ country of birth and AI personality on ad quality measures (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.
Ad performance
We also evaluated the real world performance of the ads produced by our human-AI teams in a large-scale field experiment on X. Our field study evaluated traditional marketing metrics, including click-through rates (CTR) and cost-per-click (CPC), as well as downstream ad behaviors, including view-through rates (VTR) and view-through duration (VTD), using a unique DocSend link for each ad (Figure 5). VTR and VTD were measured to assess whether participants created click-bait ads irrelevant to the research report, which could elevate CTR but reduce VTR and VTD. In line with previous findings (ju2025collaborating), our mixed effects regression analysis revealed that higher quality ad text reduced CPC while higher quality images improved CTRs.
Figure 5: Regression coefficients for human-AI personality interaction terms and ad quality ratings on click-through rates (CTR; basis points) and cost-per-click (CPC; \$) and view-through rates (VTR; %) (p<<0.05). Higher values indicate better performance for CTR and VTR, whereas lower values indicate better performance for CPC. Error bars show 95% confidence intervals. See SI Section K for the full regression tables.
Personality pairings also significantly influenced ad performance. For conscientious humans, neurotic AI lowered CTRs, while for neurotic humans, neurotic AI increased CTR performance. Conscientious humans with neurotic AI and extroverted humans with agreeable AI both experienced significantly higher CPC. Open humans experienced lower VTRs when paired with conscientious AI, while conscientious humans produced ads with higher VTRs when paired with open AI. We found no statistically significant effects for VTD. Together, these results indicate that human-AI personality pairings influence ad performance both by shaping quality metrics and through effects on ad performance not captured by quality measures.
Productivity
In addition to evaluating the quality of the ads produced, we examined the effects of personality pairings on the productivity of human-AI teams. We defined productivity as the number of ad submissions completed by participants within a 40-minute task period. Since participants were given explicit monetary incentives to produce a greater number of ads and to produce higher quality ads, this definition offers a quantitative measure of output, and thus productivity, in the collaborative process of creating display ads.
Our analysis revealed several significant effects of personality pairing on productivity (Figure 6, see SI Section H). For extroverted humans, agreeable AI agents increased productivity, and for agreeable humans, open AI increased productivity while neurotic AI decreased productivity. These findings indicate that specific personality pairings influence productivity in human-AI collaboration. In particular, neuroticism mediates how agreeable, or sycophantic, AI impacts productivity. There were no significant average treatment effects for AI personality alone (see SI Section J).
Figure 6: Regression coefficients for human-AI personality interaction terms for productivity (i.e., the number of submissions; p<<0.05) for three human personality covariates, extroversion, agreeableness, and neuroticism. Error bars show 95% confidence intervals. See SI Section H for the full regression table.
Interestingly, some personality pairings introduced a trade-off between quality and productivity. Pairing agreeable humans with open AI increased productivity but decreased image quality. Pairing agreeable humans with neurotic AI had the opposite effect—they produced fewer ads but with higher text quality. These results suggest a trade-off in human-AI collaboration whereby some personality pairings improve the productivity of human-AI teams in terms of the volume of output produced per unit time, while other pairings improve ad quality while degrading productivity.
Discussion
We provide the first experimental evidence that personality alignment improves human-AI collaboration and team performance. Our work highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success. Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance (krakowski2025human). To our knowledge, ours is the first study to measure the productivity, performance and teamwork effects of AI personalization in a context in which AI personalities are directly and randomly manipulated. As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of productivity and performance in real agentic workflows and unlock more effective and satisfying collaborations.
Our results extend the person-team fit literature to human-AI teams and reveal key patterns in human-AI collaboration dynamics (barrick1998gma, bell2007meta). For example, conscientious human-AI pairs demonstrated higher effort and cohesion, suggesting that shared conscientiousness creates a structured collaborative environment that promotes more effective interaction. In addition, agreeable AI, which is in essence sycophantic AI (sharma2025sycophancy), increased productivity for extroverted and neurotic people, but reduced all quality ratings when paired with conscientious people. This suggests that sycophantic or agreeable AI enables certain humans to produce more output per unit time, but erodes the quality of that output, perhaps because it pushes back less against human work and therefore fails to improve the work while letting a greater amount of lower quality work through to production. While agreeable AI can lead to over-reliance on AI in psychosocial contexts (fang2025psychosocial), our results suggest they can also be tuned to shape performance and productivity when tailored for individual human characteristics. This suggests that sycophantic AI could be managed toward productive outcomes when applied in the right contexts for the right people and avoided in the wrong contexts and for the wrong people. These findings highlight how different personality alignments activate distinct collaborative mechanisms, with some improving teamwork and others increasing productivity. The variation in productive outcomes associated with personality matching further demonstrates that effective human-AI collaboration requires strategic consideration of trait compatibility rather than a one-size-fits-all approach.
To make the practical implications of these results concrete, consider several vignettes that suggest particular pairing choices for conscientious, agreeable, extroverted and neurotic human workers. On one hand, conscientious people benefit from conscientious and open AI. On the other hand, agreeable people are more productive but produce lower quality work when working with open AI and produce higher quality work at a slower rate when working with neurotic AI. Extroverts benefit from extroverted AI, but produce lower quality work when collaborating with conscientious AI. Neurotic people perform better when paired with extroverted AI, but worse when working with conscientious AI. These vignettes highlight concrete design choices for matching AI personalities to workers to improve human-AI collaboration.
Although our study is the first randomized control trial of AI personalization in a real-world work task and provides the first empirical evidence that human-AI personality alignment can improve collaboration, productivity and performance in human-AI teams, it is not without limitations. First, the exact personality pairing effects discovered in our marketing context may not extend to other collaborative tasks, such as coding or analytics. Indeed, our analysis of the productivity and performance effects of human-AI personality alignment in a multi-modal ads creation task, which involved both visual ad images and textual ad copy, itself revealed that personality pairing effects can vary across tasks. Future research is therefore needed to measure pairing effects and their generalizability or jaggedness across tasks.
Second, our personality prompts were independently randomized and analyzed as interaction effects with human personalities. As such, we did not examine the performance effects of any optimization procedure or intentional pairing of human to AI personality traits. In this way, our results provide a rigorous, large-scale existence proof for the potential of optimized AI personalization, rather than exploring hypothesis driven, intentional or strategic personalization per se. While our extensive randomization allows us to analyze a wide range of diverse effects to efficiently generate hypotheses, this approach requires a large sample to have sufficient statistical power and becomes challenging to analyze in a theoretically driven manner. Future studies could focus on specific personality pairings to directly and efficiently answer theoretical hypotheses or evaluate optimized personalization, especially given the hypotheses generated from this study. Furthermore, while we employ a version of P2 prompting (jiang2023evaluating), new and developing methods may better steer AI toward desired personalities (chen2025persona), and these methods should also be investigated as avenues toward strategic AI personalization.
Third, our focus on the Big Five personality traits, though robust and widely used in human personality research, may overlook other personality traits potentially more relevant to human-AI dynamics. Our analyses indicate that demographics, like country of origin, moderate performance effects, suggesting broader heterogeneity in trait interactions that warrant further exploration. Future work could also investigate linguistic patterns to better understand traits that influence teamwork (wan2025personalized).
Fourth, the study’s single 40-minute session also limits understanding of how these relationships develop over time, such as over-reliance on AI (fang2025psychosocial). Overall, these considerations highlight avenues for future research, including in-depth textual and interaction analyses, to expand on this study’s contributions to human-AI collaboration.
Finally, our study says little about the psychosocial implications of AI personalization. Although our teamwork variables address self-assessed teamwork dynamics, we examine no measures of well-being or outcomes related to, for example, AI psychosis. As the field of AI personalization develops, it will be important to understand the psychological, social, and psychosocial implications of tailoring AI personalities to their human users and we encourage the development of that line work in future research.
Taken together, the effects shown here underscore the critical role of personality alignment in improving human-AI collaboration. By demonstrating that specific pairings can improve performance while others hinder it, we highlight the nuanced dynamics at play in human-AI teams. The observed trade-offs between quality and productivity further illustrate the complexity of these interactions. These insights pave the way for designing AI agents that are not only technically proficient but also socially and behaviorally adaptive, ensuring they complement their human counterparts effectively. Ultimately, our work demonstrates that personality-aligned AI has the potential to transform collaborative workflows and performance in human-AI teams.
Methods
This research was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects (see osf.io/jfzha and osf.io/95dhu).
The Pairit platform
The study was conducted on the Pairit platform (Figure 7). On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E, and write copy for the display ads. In the human-AI teams, the AI has full context of all elements on the platform, including screenshots of the ad images as they evolve over time during the collaboration. In all conditions, only the participants can submit ads. Additional details on the platform are in SI Section A-B.
Figure 7: The Pairit platform. The task workspace is on the left, and the messaging interface on the right. The participant chats with an AI agent which has full context of the user interface and can modify all elements on the task panel.
To enable the AI agents to act as effective collaborators on the Pairit platform, we provided them with full context of the task and collaboration by prompting each LLM call with all information on the user interface. The prompts included the same text of the task given to participants, previous submissions, personality prompts, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, general instructions, and a screenshot of the image after each change by the participant. We used OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06). Additional details in the AI agent are in SI Section  D.
Using Pairit, we designed the AI agent to be able to take the same actions that a human participant could take, apart from submitting ads. The actions include sending chat messages, editing any element of the ad copy (including the headline, primary text, and description), selecting images, generating images using the Dall-E 3 interface, or waiting (i.e., not taking any action). The agent is prompted every 10 seconds on whether to engage to take an action.
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (wei2023chainofthought). Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message.
For the randomization of AI personalities, we prompt the model with prompts that induce high or low levels of the Big Five personality traits. Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (jiang2023evaluating). See SI Section E for covariate balance of the randomization of each personality prompt.
Procedure
As soon as participants were redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 1A). In the human-human condition, which we do not discuss in this study, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, which is the focus of this study, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey. Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (rammstedt2007personality). See SI Section E for summary statistics and covariate balance of the human-AI randomization and SI Section G for the survey items.
Following the pairing and completion of the pre-task survey, participants entered a collaborative workspace where they could chat and collaboratively develop and submit display ads (Figure 1B). Participants were allotted 40 minutes to create and submit as many display ads as possible. After 40 minutes, participants were automatically re-directed to a post-task survey.
Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (hoegl2001). The original instrument included 38 questions on a 7-point Likert scale, but three communication-related items were excluded as they were irrelevant to a one-time online collaborative effort. Next, the survey included two items on prior AI experience (i.e., “I have previously used AI chatbots (e.g., ChatGPT, Bard),” and “My experiences with AI chatbots have been positive”), one item on whether they perceived their partner as AI (i.e., “I thought my collaborator was an AI during the task”), and a final item disclosing the partner’s identity (human or AI) and asking if this altered their view of the collaboration quality (i.e., “Your collaborator was an AI/a human. How has this knowledge affected your perception of the collaboration’s quality?”). 85% of participants completed the post-task survey.
Participants
We pre-registered a target sample of 2,500 individuals recruited via Prolific (prolific.com) located in the US, with balanced representation by gender and ethnicity. Ultimately, 2310 participants completed the task. Of the 2500 individuals who entered, 23 failed to join the queue and were excluded from the study. Additionally, 167 participants either withdrew or exceeded the time limit prior to being paired with a collaborator. The overall attrition rate was 7.6%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
Incentives
To incentivize the production of high-quality ads, we informed the participants of their eligibility for additional \$100 as bonus payment, determined by the volume and quality of their submissions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
Human evaluation of ad quality
To obtain human ratings of ad quality, we recruited a separate sample of 1,995 participants from Prolific based in the US, with representative stratification across gender, age, and ethnicity. A total of 1,200 individuals entered the survey. Of these participants, 5 dropped out before submitting their surveys, resulting in a dropout rate of 0.42%. We built a custom platform for this survey, run on Google Cloud Platform’s App Engine. The code is available on GitHub. This evaluation was conducted from November 7 to 9, 2024, and took a median of 16.47 minutes to complete.
To obtain ratings for all ads while avoiding survey fatigue, we created a random sample, with random order, of 40 ads per participant. To ensure that each ad received at least 3 ratings, we produced a set of 1,300 samples. As a participant entered our survey platform, we drew one sample from the set, without replacement, to provide to the participant, with each participant receiving a unique sample.
To estimate ad performance, we created and used mockups of display ads in the survey (see SI Section C). For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
Field evaluation of ad performance
We evaluated real-world ad performance on the X platform, pre-registered at osf.io/95dhu. Key metrics included click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; fraction of document viewed), and view-through duration (VTD; seconds). VTR and VTD were tracked via a unique DocSend link for each ad. Our ad campaigns assessed causal ad effects conditional on online algorithms and divergent targeting (braun2024adplatforms, braun2024abtesting). No holdout group was used, as unique DocSend links imply zero visits without ads (braun2024adplatforms).
We sampled 2,000 ads from 11,138, selecting 1–2 ads per 1,834 human-human and human-AI teams using stratified sampling. Ads were grouped by team type and divided into 10 strata based on human-predicted click likelihood. Eight ads were excluded for potential policy violations (e.g., violence, sexual, or drug content). To avoid audience overlap, 400 campaigns each ran five unique ads as split tests. Each campaign targeted 133 random ZIP codes with population between 10,000 and 100,000, as per the 2020 American Community Survey. One-way ANOVA yielded non-significant results across ZIP codes: $F(399,53199)=0.954$, $p=0.734$ for population; $F(399,53199)=0.973$, $p=0.636$ for income. Ads were run from January 21 to February 9, 2025, with 50 campaigns active for two days each due to platform constraints. Campaign random effects controlled for temporal confounders. In our regression analyses, advertising spend was controlled for to account for auto-bidding.
Model specifications
To evaluate the effect of personality trait interactions on human-AI team outcomes, we use the following regression model:
$Y_{J}=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}+\epsilon_{i}\text{,}$ (1)
where $Y_{J}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low), and $\epsilon_{i}$ is the error term. The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating (with standard errors clustered at the ad level); for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors). The outcomes of interest $Y_{i}$ include human-evaluated ad quality, productivity (i.e., the number of ads submitted), and teamwork quality measures.
To evaluate the field experiment, we use the following mixed effects model:
$Y_{ic}=$ $\ \sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}$
$+\text{Text}_{ic}+\text{Image}_{ic}+\text{Click}_{ic}+\text{Spend}_{ic}+u_{c}+\epsilon_{ic}\text{,}$ (2)
where $Y_{ic}$ represents the outcome (e.g., CPC, CTR, VTR, VTD) for ad $i$ in campaign $c$, Image, Text, and Click are human-rated quality measures, $\text{Spend}_{ic}$ is the campaign spend, $u_{c}\sim N(0,\sigma^{2}_{u})$ is the random intercept for campaign $c$, and the rest are the same as in Equation 1.
Acknowledgments
We thank Dean Eckles, John Horton and members of the MIT Social Analytics Lab for their invaluable discussions. This work was supported by the MIT Initiative on the Digital Economy (IDE) at the MIT Sloan School of Management and was approved by the Massachusetts Institute of Technology Institutional Review Board (IRB).

### v2

@@
 Abstract
-Here we ask how AI agent “personalities” interact with human personalities, and other traits, to shape human-AI collaboration, productivity and performance.
-To estimate these relationships, we conducted a large-scale preregistered randomized experiment that paired 1,258 participants with AI agents that were prompted to exhibit varying levels of the Big Five personality traits.
-These human-AI teams produced 7,266 display ads for a real think tank, and the quality of these ads was evaluated by 1,995 independent human raters as well as in a field experiment conducted on X, which generated nearly 5 million impressions.
-We found, first, that personality pairing impacted teamwork quality.
-For example, neurotic AI improved teamwork for agreeable humans but impaired it for conscientious humans.
-Second, we found productivity effects of personality pairing and a “productivity-performance trade-off" in which certain pairings (e.g., agreeable human with neurotic AI) produced fewer ads but of higher quality.
-Third, personality pairing influenced ad quality and performance.
-For example, quality improved when open humans were paired with conscientious AI and when conscientious humans were paired with disagreeable AI.
-Some of these pairing effects were “jagged” in that they varied across text and visual tasks.
-For example open humans produced higher quality images but lower quality text when paired with agreeable AI.
-Pairing effects were also present in other human traits, like country of origin.
-For example, extroverted AI improved quality for Latin American workers, but degraded quality for East Asian workers.
-These findings demonstrate that human-AI personality alignment significantly improves collaboration, productivity, and performance and lay a foundation for future research on improving human-AI collaboration through AI personalization.
-Introduction
+Here we examine how AI agent “personalities” interact with human personalities to shape human-AI collaboration and performance.
+In a large-scale, preregistered randomized experiment, we paired 1,258 participants with AI agents prompted to exhibit varying levels of the Big Five personality traits.
+These human-AI teams produced 7,266 display ads for a real think tank, which we evaluated using 1,995 independent human raters and a field experiment on X that generated nearly 5 million impressions.
+We found that human and AI personalities individually shaped ad quality and teamwork.
+When examined together, human-AI personality pairings directly effected ad quality outcomes.
+For example, extraverted humans paired with conscientious AI produced the lowest-quality ads, followed by conscientious humans paired with agreeable AI and neurotic humans paired with conscientious AI.
+In the field experiment, ad quality significantly influenced ad performance, measured by click-through rates and cost-per-click, and neurotic humans paired with neurotic AI achieved higher click-through rates, even after controlling for ad quality.
+Together, these results provide the first large-scale causal experimental evidence that specific personality pairings can improve human-AI collaboration and motivate future research on the implications of AI personalization for performance and teamwork dynamics in human-AI teams.
 Artificial Intelligence (AI) tools have gained significant attention for their ability to increase productivity and performance across a wide range of tasks.
-For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (noy2023experimental).
-AI-enhanced resumes increased job hiring rates by 8% (wiles2023writing).
-In customer support settings, AI assistance increased productivity by an average of 14% (brynjolfsson2023genai).
-But these benefits also exhibited considerable heterogeneity, varying across tasks (dellacqua2023navigating, ju2025collaborating), worker seniority (wang2023friend), gender (dellacqua2023navigating, humlum2024adoption), and skill levels (noy2023experimental, brynjolfsson2023genai, otis2024uneven).
-In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (chen2024impact, yu2024heterogeneity, agarwal2023combining).
-In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (yang2025my, vaccaro2024combinations, sergeyuk2025human).
+For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (Noy and Zhang, 2023).
+AI-enhanced resumes increased job hiring rates by 8% (Wiles et al., 2023).
+In customer support settings, AI assistance increased productivity by an average of 14% (Brynjolfsson et al., 2023).
+But these benefits also exhibited considerable heterogeneity, varying across tasks (Dell’Acqua et al., 2023, Ju and Aral, 2025), worker seniority (Wang et al., 2023), gender (Dell’Acqua et al., 2023, Humlum and Vestergaard, 2024), and skill levels (Noy and Zhang, 2023, Brynjolfsson et al., 2023, Otis et al., 2024).
+In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (Chen et al., 2024, Yu et al., 2024, Agarwal et al., 2023).
+In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (Yang et al., 2025, Vaccaro et al., 2024, Sergeyuk et al., 2025).
 Unfortunately, experiments documenting these productivity and performance gains have generally provided participants with identical AI tools, leaving open the possibility that at least some of the observed variation in performance stems from a mismatch between AI and individual human characteristics.
@@
 Such personalization may now be possible as foundation model companies have recently begun developing, refining and deploying AI personalities to improve human-AI interaction.
-For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities (openai2025customizing, openai2025sycophancy).
-Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences (anthropic2024character, sharma2025sycophancy), or other “evil” traits, and introduced novel training methods to mitigate such personality shifts in their models.
+For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities OpenAI (2025b, a).
+Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences Anthropic (2024), Sharma et al.
+(2025), and introduced novel training methods to mitigate such personality shifts in their models.
 These advances enable AI agents to be tailored to users, potentially improving human-AI collaboration and ushering in a new era of AI personalization.
 Despite the development of AI personalities, however, no prior research has rigorously and experimentally examined their causal effects on human-AI collaboration.
-Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork, productivity, and performance as the lack experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, AI psychoses.
-In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (zhang2023teammate, hemmer2023, woolley2023collective).
-Humans exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (kichuk1997, peeters2006, woolley2010ci).
-Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (barrick1998gma, bell2007meta).
-As AI has evolved from a passive tool to an active collaborator (makarius2020rising, fugener2022cognitive, anthony2023, collins2024building, ju2025collaborating), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time.
-Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (li2025acceptance, tully2025lower, fang2025psychosocial).
-Unfortunately, although human and AI traits shape human-AI teamwork, their interaction remains understudied (krakowski2025human).
-With AI agents increasingly operating as social actors in collaborative workflows (anthony2023, gelbrich2025), it is crucial to understand whether the principles of personality alignment apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts.
-To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork, productivity, and performance.
-Our study leverages the unique capabilities of AI agents to vary their individual personalities (jiang2023evaluating, chen2025persona), allowing us to systematically investigate how human-AI personality alignment influences the productivity and performance of human-AI teams.
-Systematic prompt variation can effectively shape AI agent behavior (schulhoff2024prompt), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (brown2020language).
-Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (wei2023chainofthought) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (chen2025persona).
+Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork and performance as the lack of experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, adverse psychosocial outcomes (Fang et al., 2025).
+In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (Zhang et al., 2023, Hemmer et al., 2023, Woolley et al., 2023).
+People exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (Kichuk and Wiesner, 1997, Peeters et al., 2006, Woolley et al., 2010).
+Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (Barrick et al., 1998, Bell, 2007).
+As AI has evolved from a passive tool to an active collaborator (Makarius et al., 2020, Fügener et al., 2022, Anthony et al., 2023, Collins et al., 2024, Ju and Aral, 2025), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time.
+Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (Li et al., 2025, Tully et al., 2025, Fang et al., 2025).
+Early work on computers as social actors showed that people prefer personality-matched computers (Nass and Moon, 2000), and personality-targeted advertising can alter behavior at scale (Matz et al., 2017, 2024).
+However, these studies examined either preference outcomes with pre-LLM systems or one-shot content personalization, not interactive collaboration with AI agents.
+The interaction between human and AI traits in collaborative production remains understudied Krakowski et al.
+(2025).
+With AI agents increasingly operating as social actors in collaborative workflows (Anthony et al., 2023, Gelbrich et al., 2025), it is crucial to understand whether the principles of personality pairing apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts.
+To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork and performance.
+Our study leverages the unique capabilities of AI agents to vary their individual behavioral styles (Jiang et al., 2023, Chen et al., 2025), allowing us to investigate how human-AI personality pairing influences the performance of human-AI teams.
+Systematic prompt variation can effectively shape AI agent behavior (Schulhoff et al., 2024), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (Brown et al., 2020).
+Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (Wei et al., 2022) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (Chen et al., 2025).
 These methods enable us to intentionally design AI agents with particular social and behavioral characteristics, which can be tailored to complement human collaborators.
-Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (jakesh2023opinionated).
+Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (Jakesch et al., 2023).
 Investigating human-AI collaboration with agentic tools is challenging in practice.
@@
 Experiment design
-To investigate how personality alignment between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S.
-population using Pairit, an experimental platform we developed for extensible human-AI collaboration (ju2025collaborating).
-After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (jiang2023evaluating).
-These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank within a 40-minute session, using real-time chat and synchronized text and image editing tools.
+To investigate how personality pairing between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S.
+population using Pairit, an experimental platform we developed for extensible human-AI collaboration (Ju and Aral, 2025).
+After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (Jiang et al., 2023).
+These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank during a 40-minute session, using real-time chat and synchronized text and image editing tools.
 Teams were provided with image assets by the think tank and could separately call Dall-E 3 to generate new images for the campaigns.
-As described in the Materials and Methods, the AI agents were fully multi-modal, had access to the Internet, could manipulate the images, the image carousel and the text of the ads, text with their human collaborators and could also call Dall-E 3 to generate new images.
-This setup allowed us to assess the causal effects of personality alignment on collaboration outcomes.
-The randomization of AI personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration, productivity and performance of human-AI teams.
+Throughout the collaboration, the AI agent dynamically received screenshots of the current ad image alongside the task brief, the current ad copy, the chat history with the human, and its own action history.
+As it was provided this context, it could choose from five action types: Wait, Chat, EditText (revise the headline, primary text, description, or image prompt), SelectImage (choose an image from the carousel), or GenerateImage (call Dall-E 3 to produce a new image).
+The agent and the human shared a single live workspace, so the human could see and edit the same ad in real time as the agent acted.
+Only the human partner could submit ads, and teams iterated to submit as many ads as they could within the 40-minute window.
+This setup allowed us to assess the causal effects of personality pairing on collaboration outcomes.
+The randomization of AI agent personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration and performance of human-AI teams.
 Figure 1: The experimental design for investigating human-AI collaboration.
@@
 (C) Pairs collaborate in a real-time collaborative workspace to create display ads.
-Importantly, we approach personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (chhikara2025mem0, sharma2025sycophancy).
-Instead, we randomize AI personalities to learn the causal effects of various personality pairings on human-AI team performance.
-This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (jung2013backchanneling, jung2018theories).
+Importantly, we approached personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (Chhikara et al., 2025, Sharma et al., 2025).
+Instead, we treated AI personality as an experimental treatment: each AI agent was independently and randomly assigned a personality profile (high or low on each Big Five personality trait), and we estimated how this assignment interacted with the human partner’s measured traits to shape collaboration outcomes.
+Random assignment rules out selection effects, such as users gravitating toward AI styles they prefer, so observed differences across pairings reflect causal effects of the pairing itself.
+This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (Jung et al., 2013, Jung and Hinds, 2018).
 To assess the quality of the display ads produced by the human-AI teams, we conducted a separate evaluation study using a second independent sample of 1,198 new participants recruited from Prolific.
@@
 Raters assessed each ad on three dimensions using 7-point Likert scales: text quality (“The text is present, clear, relevant, and engaging”), image quality (“The image is visually appealing”), and likelihood of clicking the ad (“I am likely to click on this ad”).
-The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Supporting Information [SI] Section C).
+The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Materials and Methods).
 To evaluate the real-world performance of the display ads, we then conducted a pre-registered field experiment on the X platform.
-We sampled 2,000 ads from the total 7,266 ads using stratified sampling based on predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution.
-These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (braun2024abtesting, braun2024adplatforms).
+We sampled 2,000 ads from the 7,266 total ads produced during the experiment, using stratified random sampling based on the predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution.
+These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (Braun and Schwartz, 2024, Braun et al., 2024).
 Key performance metrics included the click-through rate (CTR) and the cost-per-click (CPC) measured on X, as well as the view-through rate (VTR; fraction of document viewed) and the view-through duration (VTD; seconds) tracked via unique DocSend links for each ad (details of the field experiment are provided in the Materials and Methods).
+We employed ZIP-code-level randomization across campaigns to prevent audience spillovers; campaign random effects were specified to absorb campaign-level variation in audience composition and temporal confounders; and controls for advertising spend accounted for any remaining differences in budget allocations across campaigns and ads created by auto-bidding.
+We applied Benjamini-Hochberg (BH) multiple hypothesis testing corrections to all analyses in the paper and report results based on these corrections throughout.
 Results
-Our findings revealed that personality pairings significantly affected collaborative outcomes including teamwork quality, productivity, ad quality and ad performance.
-For example, humans high in conscientiousness created ad text of greater quality when paired with conscientious AI agents, while extroverted humans paired with conscientious AI produced lower quality text, lower quality images, and ads with lower click likelihood.
-Some pairings demonstrated uneven “jagged” performance patterns, with effects varying sharply between the textual and visual components of the task.
-We further observed effects on productivity, with agreeable humans, for example, producing more ads when paired with open AI.
-Additionally, certain combinations yielded a performance-productivity trade-off, such as agreeable humans with neurotic AI generating fewer ads, but ads of higher quality.
-Overall, these results underscore the substantial role of personality alignment in human-AI teamwork and lay a foundation for future research on AI personalization to improve human-AI productivity and collaborative performance.
-In particular, we found specific effects on teamwork, productivity and output quality.
-Teamwork
-First, we found personality pairings significantly influenced teamwork quality.
-To evaluate these effects, participants completed a 35-item teamwork quality survey after the task, which assessed six facets of collaboration: communication, coordination, balance of member contributions, mutual support, effort, and cohesion (hoegl2001).
-We conducted a regression analysis on data produced by our experiments to identify the average treatment effects of AI personalities as well as the interaction effects of human and AI personality traits in combination and found statistically significant effects for human openness, conscientiousness, and agreeableness, as well as for AI neuroticism (Figure 2, see SI Section H).
-Figure 2: Regression coefficients for human-AI personality interaction terms for teamwork quality (p<<0.05) (hoegl2001).
+Our results demonstrate that personality pairing affected ad quality, teamwork, and real-world ad performance in our experiments.
+We first report the individual effects of human and AI personality traits on teamwork and ad quality, then turn to the personality pairing results and conclude by reporting the results of the field experiment.
+Individual trait effects on teamwork and ad quality
+We first tested whether measured human traits and randomized AI traits independently predicted teamwork and ad quality in human-AI teams (Figure 2).
+Each model regressed an outcome on the five human trait scores and the five AI trait scores with no interaction terms, and BH corrections were applied within each outcome family.
+Since AI traits were experimentally randomized, while human traits were measured but not manipulated, the AI coefficients can be interpreted causally, whereas the human trait coefficients represent correlations not causal relationships.
+Figure 2: Effect of personality on ad quality and teamwork.
+Coefficients from regressions of each outcome on the five human trait scores and the five AI trait scores, with no interaction terms, BH-corrected within each outcome family.
+Panel A shows ad quality dimensions (text, image, click likelihood); panel B shows teamwork indices (task, social).
+Within each panel, human trait effects appear on the left and AI trait effects on the right.
+X-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism.
+Symbols indicate significance after BH correction: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
 Error bars show 95% confidence intervals.
-See SI Section H for the full regression tables.
-In examining interaction effects, we found that for open humans, conscientious AI lowered the balance of member contributions, mutual support, effort, and cohesion.
-For conscientious humans, open AI improved balance and cohesion, and conscientious AI improved effort and cohesion.
-However, for conscientious humans, neurotic AI lowered communication, coordination, balance, and mutual support.
-Finally, for agreeable humans, neurotic AI improved communication, coordination, mutual support, and effort.
-However, these positive results reversed when agreeable humans were paired with open AI (for cohesion) and with conscientious AI (for coordination, balance, and cohesion).
-These effects revealed that personality pairings shape perceived teamwork dynamics in human-AI collaboration.
-Ad quality
-In human evaluations of ad quality, we found that personality pairing between humans and AI agents significantly influenced ad text quality (Figure 3) beyond the effects of AI personality alone (see SI Section H).
-The average treatment effects demonstrated that conscientious AI reduced, and extroverted AI improved, text quality on average.
-The interaction effects then demonstrated statistically and economically significant effects of human-AI personality pairing on ad quality.
-For both open and conscientious humans, conscientious AI increased text quality, but agreeable AI decreased text quality.
-For extroverted humans, extroverted AI increased text quality, but conscientious AI decreased text quality.
-For agreeable humans, neurotic AI increased text quality, but extroverted AI decreased text quality.
-These results indicate that aligning AI personality traits with human traits impacts the effectiveness of collaborative ad text production.
-Figure 3: Regression coefficients for human-AI personality interaction terms on ad text quality, image quality, and click likelihood (p<<0.05).
+See SI Appendix, Section LABEL:si:full, for the full regression tables.
+Human traits predicted both teamwork and ad quality.
+Conscientious and agreeable humans experienced stronger task ($p<0.001$, $p<0.01$) and social teamwork ($p<0.001$, $p<0.001$), while neurotic humans reported weaker task teamwork ($p<0.05$).
+Conscientiousness and agreeableness in humans also predicted higher ratings on all six underlying teamwork dimensions (communication, coordination, balance of contributions, mutual support, effort, and cohesion) (see SI Appendix, Section LABEL:si:main-effects).
+Human openness predicted higher text quality ($p<0.05$) and human extraversion predicted lower text quality ($p<0.05$).
+AI traits also affected ad quality: AI extraversion increased text quality ($p<0.05$), whereas AI conscientiousness reduced text quality ($p<0.01$) (see SI Appendix, Section LABEL:si:main-effects, for the full regression tables).
+It is perhaps not surprising that human and AI personalities affect teamwork and ad quality.
+But evaluating the potential benefits of AI personalization requires a more comprehensive analysis of human-AI personality pairings.
+Figure 3: Effect of personality pairing on ad quality.
+Coefficients for human-AI personality interaction terms on the composite ad quality index (gray) and the three ad quality dimensions: text quality (blue), image quality (red), and click likelihood (green), BH-corrected within each outcome family.
+Panel titles show the full human trait names; x-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism.
+Symbols indicate significance after BH correction: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
 Error bars show 95% confidence intervals.
-See SI Section H for the full regression table.
-For image quality, we observed a different set of significant interaction effects based on human-AI personality pairing.
-For open humans, agreeable AI increased image quality while, for conscientious humans, open AI increased image quality.
-For extroverts and agreeable participants, being paired with conscientious and open AI, respectively, decreased image quality.
-For click likelihood ratings, which reflect the perceived engagement potential of the ads, we observed no significant effects for humans with high openness or agreeableness.
-For conscientious humans, agreeable AI decreased click likelihood, while for extroverted and neurotic humans, conscientious AI decreased click likelihood.
-Across the three measures of ad quality (text quality, image quality and perceived click likelihood), personality pairings between humans and AI agents consistently influenced outcomes.
-For extroverts, conscientious AI decreased quality across all measures.
-For conscientious humans, agreeable AI decreased text quality and click likelihood.
-Some pairings had significant and opposite effects, indicating a “jagged edge” of human-AI competence across tasks and personality pairings (dellacqua2023navigating).
-For example, for open humans, agreeable AI improved image quality but worsened text quality.
-Collectively, these patterns suggest that strategic personality pairings can optimize human-AI collaboration outcomes, with some generalizable and some task-specific personality pairings.
-In addition to participant personality traits, we found that participant demographics, including previous AI experience, employment status, and country of origin, moderated the effects of AI personality traits on ad quality.
-For individuals with previous AI experience, extroverted AI hurt overall performance, while neurotic AI improved performance (see SI Sections G).
-Although all participants were based in and representative of the demographic composition of the U.S., their diverse countries of birth allowed us to explore possible cultural and regional differences in collaborative outcomes (Figure 4, see SI Section H).
-Participants from Western regions, the Middle East & South Asia produced significantly lower quality images when paired with agreeable AI, while participants from East Asia produced images of significantly lower quality when paired with extroverted and agreeable AI.
-Participants from Latin America produced significantly higher quality ad text when paired with extroverted AI and significantly lower quality ad text when paired with conscientious, agreeable, and neurotic AI.
-For participants from all regions except East Asia, agreeable AI significantly lowered click likelihood.
-For those from Latin America, extroverted AI improved click likelihood, while conscientious AI lowered click likelihood.
-For those from the Middle East & South Asia, agreeable AI lowered click likelihood.
-Interestingly, we observed that agreeable AI consistently undermined performance across all measures, reducing image quality (for Western, East Asian, Middle Eastern & South Asian workers), text quality (for Latin American workers), and click likelihood (for Western, Latin American, Middle Eastern & South Asian workers).
-Taken together, these patterns suggest individual characteristics like culture mediate how AI personality traits, especially agreeable or sycophantic AI (sharma2025sycophancy), affect collaboration and suggest the potential for future research investigating the strategic pairing of AI personalities with various human traits.
-Figure 4: Regression coefficients for interaction terms of a participants’ country of birth and AI personality on ad quality measures (p<<0.05).
+See SI Appendix, Section LABEL:si:full, for the full regression table.
+Personality pairing and ad quality
+Beyond individual traits, personality pairings between humans and AI agents significantly influenced ad quality (preregistered; Figure 3; see SI Appendix, Section LABEL:si:prereg).
+Three pairings drove the strongest effects on the composite ad quality index and all three reduced quality.
+Extraverted humans paired with conscientious AI produced the lowest-quality ads ($p<0.001$), followed by conscientious humans paired with agreeable AI ($p=0.01$) and neurotic humans paired with conscientious AI ($p=0.049$).
+These results suggest a broad pattern in personality pairing effects on human-AI collaboration: when humans who are assertive, goal driven and set in their ways (traits associated with extraversion and conscientiousness) collaborate with compliant AI that are not skeptical of and do not push back on human ideas, output quality decreases.
+For example, when extraverted humans, who are assertive, enthusiastic and “high energy," collaborate with conscientious AI, the result is a collaboration that efficiently and effectively implements the human’s plans, with little explicit push back from AI on whether they are the right plans.
+When conscientious humans, who are disciplined and goal-directed, collaborate with agreeable AI, the result is a more compliant implementation of human ideas.
+When neurotic humans, who are more reactive to stress and uncertainty, collaborate with conscientious AI, the team implements a more volatile and less stable collaborative process, again with limited explicit push back from the AI, resulting in lower quality output.
+Exploratory LLM labeling of the human-AI chat logs is consistent with this interpretation.
+Relative to the corresponding low-trait AI comparison cells, conscientious AI paired with extraverted humans showed slightly less AI questioning (0.331 vs.
+0.335) and slightly more supportiveness (0.382 vs.
+0.376), while the humans themselves were slightly more assertive (0.213 vs.
+0.202).
+Agreeable AI paired with conscientious humans showed lower AI push back (0.003 vs.
+0.004) and higher order-taking (0.107 vs.
+0.104), while the humans themselves were more goal-oriented (0.239 vs.
+0.212).
+Conscientious AI paired with neurotic humans again showed lower AI push back (0.004 vs.
+0.005) and higher order-taking (0.114 vs.
+0.109) (see SI Appendix, Section LABEL:si:sycophancy).
+Overall, explicit AI push back was rare in the human-AI chats ($1\%$ of messages).
+This suggests that recent trends toward AI sycophancy are likely to produce lower quality outcomes in human-AI collaboration and that extraverted, conscientious and neurotic humans are likely to be the most vulnerable to performance degradation when collaborating with sycophantic AI.
+Although this linguistic analysis is exploratory and not necessary to confirm the conclusions of our statistical analysis, that linguistic indicators conform to the patterns in our statistical results increases our confidence in our interpretations and conclusions.
+At the level individual performance indicators, extraverted humans paired with conscientious AI lowered quality across all three measured dimensions including text ($p<.001$), image ($p<.001$), and click likelihood ($p<.05$), while pairing with extraverted AI increased text quality ($p<.01$).
+Conscientious humans produced lower quality text when paired with agreeable AI ($p<.001$) and higher quality text and images when paired with conscientious ($p<.01$) and open AI ($p<.05$).
+Open humans produced higher quality text when paired with conscientious AI ($p<.05$) and lower quality text when paired with agreeable AI ($p<.05$).
+Agreeable humans produced lower quality text when paired with extraverted AI ($p<.05$) and lower quality images when paired with open AI ($p<.05$).
+Finally, neurotic humans produced ads with lower click likelihood when paired with conscientious AI ($p<.05$).
+AI experience (preregistered) moderated how AI personality affected ad quality, while employment status and country of origin showed additional exploratory heterogeneity (see SI Appendix, Section LABEL:si:demographics; see also SI Appendix, Section LABEL:si:prereg).
+Taken together, these results suggest the existence of economically and statistically significant gains from AI personalization.
+When people collaborate with AI with personalities that are more appropriate for them, the quality of their output improves significantly.
+Personality pairing and teamwork
+Personality pairings also shaped the social side of teamwork, though more weakly than ad quality (preregistered; see SI Appendix, Section LABEL:si:prereg).
+Participants completed a 35-item teamwork quality survey after the task (Hoegl and Gemuenden, 2001), which assessed six facets of teamwork grouped into a social dimension (mutual support, effort, and cohesion) and a task dimension (communication, coordination, and balance of member contributions).
+The social dimension showed joint significance of personality pairing effects ($F$-test $p=0.022$; MANOVA $p=0.004$), although no individual interactions survived the BH corrections, and the task dimension did not show joint significance.
+These results suggest that personality pairings may influence collaboration dynamics, though individual effects require confirmation in future targeted experiments.
+See SI Appendix, Section LABEL:si:full, for the full coefficient plots and regression tables.
+Ad performance
+Having established the effects of personality pairing on human and AI rated ad quality, we evaluated the real-world performance of the ads in a large-scale field experiment on X (Braun et al., 2024, Braun and Schwartz, 2024).
+The field experiment measured click-through rates (CTR), cost-per-click (CPC), view-through rates (VTR), and view-through duration (VTD) using unique DocSend links for each ad.
+The quality of the ads predicted their field performance as higher quality ad text reduced the cost per click (CPC) of the ads, while higher quality images increased the click through rate (CTR) on the ads (Figure 4A).
+These results trace the impact of personalization on performance, as personality pairings affect ad quality and ad quality directly affects ad performance in the field.
+Figure 4: Effect of personality pairing on field performance.
+Mixed-effects regression coefficients with campaign random effects.
+Panel A shows the effect of ad quality (image, text, click likelihood) on click-through rate (CTR) and cost-per-click (CPC); symbols reflect raw $p$-values from the quality-only regression.
+Panel B shows human-AI personality interaction coefficients on CTR after controlling for ad quality (image, text, click, spend); subplot titles show the full human trait names and x-axis labels use O = Openness, C = Conscientiousness, E = Extraversion, A = Agreeableness, and N = Neuroticism.
+Panel B symbols indicate significance after BH correction within the 25-interaction CTR testing family: $^{\dagger}p<0.10$, $^{*}p<0.05$, $^{**}p<0.01$, $^{***}p<0.001$.
 Error bars show 95% confidence intervals.
-See SI Section H for the full regression table.
-Ad performance
-We also evaluated the real world performance of the ads produced by our human-AI teams in a large-scale field experiment on X.
-Our field study evaluated traditional marketing metrics, including click-through rates (CTR) and cost-per-click (CPC), as well as downstream ad behaviors, including view-through rates (VTR) and view-through duration (VTD), using a unique DocSend link for each ad (Figure 5).
-VTR and VTD were measured to assess whether participants created click-bait ads irrelevant to the research report, which could elevate CTR but reduce VTR and VTD.
-In line with previous findings (ju2025collaborating), our mixed effects regression analysis revealed that higher quality ad text reduced CPC while higher quality images improved CTRs.
-Figure 5: Regression coefficients for human-AI personality interaction terms and ad quality ratings on click-through rates (CTR; basis points) and cost-per-click (CPC; \$) and view-through rates (VTR; %) (p<<0.05).
-Higher values indicate better performance for CTR and VTR, whereas lower values indicate better performance for CPC.
-Error bars show 95% confidence intervals.
-See SI Section K for the full regression tables.
-Personality pairings also significantly influenced ad performance.
-For conscientious humans, neurotic AI lowered CTRs, while for neurotic humans, neurotic AI increased CTR performance.
-Conscientious humans with neurotic AI and extroverted humans with agreeable AI both experienced significantly higher CPC.
-Open humans experienced lower VTRs when paired with conscientious AI, while conscientious humans produced ads with higher VTRs when paired with open AI.
-We found no statistically significant effects for VTD.
-Together, these results indicate that human-AI personality pairings influence ad performance both by shaping quality metrics and through effects on ad performance not captured by quality measures.
-Productivity
-In addition to evaluating the quality of the ads produced, we examined the effects of personality pairings on the productivity of human-AI teams.
-We defined productivity as the number of ad submissions completed by participants within a 40-minute task period.
-Since participants were given explicit monetary incentives to produce a greater number of ads and to produce higher quality ads, this definition offers a quantitative measure of output, and thus productivity, in the collaborative process of creating display ads.
-Our analysis revealed several significant effects of personality pairing on productivity (Figure 6, see SI Section H).
-For extroverted humans, agreeable AI agents increased productivity, and for agreeable humans, open AI increased productivity while neurotic AI decreased productivity.
-These findings indicate that specific personality pairings influence productivity in human-AI collaboration.
-In particular, neuroticism mediates how agreeable, or sycophantic, AI impacts productivity.
-There were no significant average treatment effects for AI personality alone (see SI Section J).
-Figure 6: Regression coefficients for human-AI personality interaction terms for productivity (i.e., the number of submissions; p<<0.05) for three human personality covariates, extroversion, agreeableness, and neuroticism.
-Error bars show 95% confidence intervals.
-See SI Section H for the full regression table.
-Interestingly, some personality pairings introduced a trade-off between quality and productivity.
-Pairing agreeable humans with open AI increased productivity but decreased image quality.
-Pairing agreeable humans with neurotic AI had the opposite effect—they produced fewer ads but with higher text quality.
-These results suggest a trade-off in human-AI collaboration whereby some personality pairings improve the productivity of human-AI teams in terms of the volume of output produced per unit time, while other pairings improve ad quality while degrading productivity.
+See SI Appendix, Section LABEL:si:full, for the full field regression tables and additional outcomes (CPC, VTR, VTD), and Section LABEL:si:robustness for BH corrections across all testing families.
+Personality pairings also influenced ad performance beyond what ad quality alone could explain (preregistered; see Materials and Methods and SI Appendix, Section LABEL:si:prereg).
+Even when controlling for image quality, text quality, and click likelihood, neurotic humans paired with neurotic AI achieved significantly higher click-through rates (Figure 4B; $p=0.036$; SI Appendix, Section LABEL:si:robustness).
+Other pairings shifted costs-per-click and view-through rates, though these additional effects did not survive multiple hypothesis testing corrections.
+These results suggest that personality pairings also influence ad performance through channels not fully captured by our ad quality measures alone (SI Appendix, Section LABEL:si:full).
+Interestingly, the same neurotic-neurotic pairing was also the only pairing that significantly elevated delegation behavior and directive language in human-AI collaboration (SI Appendix, Section LABEL:si:interaction-content), a process feature shown in prior work to predict click ratings in human-AI teams (Ju and Aral, 2025).
 Discussion
-We provide the first experimental evidence that personality alignment improves human-AI collaboration and team performance.
-Our work highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success.
-Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance (krakowski2025human).
-To our knowledge, ours is the first study to measure the productivity, performance and teamwork effects of AI personalization in a context in which AI personalities are directly and randomly manipulated.
-As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of productivity and performance in real agentic workflows and unlock more effective and satisfying collaborations.
-Our results extend the person-team fit literature to human-AI teams and reveal key patterns in human-AI collaboration dynamics (barrick1998gma, bell2007meta).
-For example, conscientious human-AI pairs demonstrated higher effort and cohesion, suggesting that shared conscientiousness creates a structured collaborative environment that promotes more effective interaction.
-In addition, agreeable AI, which is in essence sycophantic AI (sharma2025sycophancy), increased productivity for extroverted and neurotic people, but reduced all quality ratings when paired with conscientious people.
-This suggests that sycophantic or agreeable AI enables certain humans to produce more output per unit time, but erodes the quality of that output, perhaps because it pushes back less against human work and therefore fails to improve the work while letting a greater amount of lower quality work through to production.
-While agreeable AI can lead to over-reliance on AI in psychosocial contexts (fang2025psychosocial), our results suggest they can also be tuned to shape performance and productivity when tailored for individual human characteristics.
-This suggests that sycophantic AI could be managed toward productive outcomes when applied in the right contexts for the right people and avoided in the wrong contexts and for the wrong people.
-These findings highlight how different personality alignments activate distinct collaborative mechanisms, with some improving teamwork and others increasing productivity.
-The variation in productive outcomes associated with personality matching further demonstrates that effective human-AI collaboration requires strategic consideration of trait compatibility rather than a one-size-fits-all approach.
-To make the practical implications of these results concrete, consider several vignettes that suggest particular pairing choices for conscientious, agreeable, extroverted and neurotic human workers.
-On one hand, conscientious people benefit from conscientious and open AI.
-On the other hand, agreeable people are more productive but produce lower quality work when working with open AI and produce higher quality work at a slower rate when working with neurotic AI.
-Extroverts benefit from extroverted AI, but produce lower quality work when collaborating with conscientious AI.
-Neurotic people perform better when paired with extroverted AI, but worse when working with conscientious AI.
-These vignettes highlight concrete design choices for matching AI personalities to workers to improve human-AI collaboration.
-Although our study is the first randomized control trial of AI personalization in a real-world work task and provides the first empirical evidence that human-AI personality alignment can improve collaboration, productivity and performance in human-AI teams, it is not without limitations.
-First, the exact personality pairing effects discovered in our marketing context may not extend to other collaborative tasks, such as coding or analytics.
-Indeed, our analysis of the productivity and performance effects of human-AI personality alignment in a multi-modal ads creation task, which involved both visual ad images and textual ad copy, itself revealed that personality pairing effects can vary across tasks.
-Future research is therefore needed to measure pairing effects and their generalizability or jaggedness across tasks.
-Second, our personality prompts were independently randomized and analyzed as interaction effects with human personalities.
-As such, we did not examine the performance effects of any optimization procedure or intentional pairing of human to AI personality traits.
-In this way, our results provide a rigorous, large-scale existence proof for the potential of optimized AI personalization, rather than exploring hypothesis driven, intentional or strategic personalization per se.
-While our extensive randomization allows us to analyze a wide range of diverse effects to efficiently generate hypotheses, this approach requires a large sample to have sufficient statistical power and becomes challenging to analyze in a theoretically driven manner.
-Future studies could focus on specific personality pairings to directly and efficiently answer theoretical hypotheses or evaluate optimized personalization, especially given the hypotheses generated from this study.
-Furthermore, while we employ a version of P2 prompting (jiang2023evaluating), new and developing methods may better steer AI toward desired personalities (chen2025persona), and these methods should also be investigated as avenues toward strategic AI personalization.
-Third, our focus on the Big Five personality traits, though robust and widely used in human personality research, may overlook other personality traits potentially more relevant to human-AI dynamics.
-Our analyses indicate that demographics, like country of origin, moderate performance effects, suggesting broader heterogeneity in trait interactions that warrant further exploration.
-Future work could also investigate linguistic patterns to better understand traits that influence teamwork (wan2025personalized).
-Fourth, the study’s single 40-minute session also limits understanding of how these relationships develop over time, such as over-reliance on AI (fang2025psychosocial).
-Overall, these considerations highlight avenues for future research, including in-depth textual and interaction analyses, to expand on this study’s contributions to human-AI collaboration.
-Finally, our study says little about the psychosocial implications of AI personalization.
-Although our teamwork variables address self-assessed teamwork dynamics, we examine no measures of well-being or outcomes related to, for example, AI psychosis.
-As the field of AI personalization develops, it will be important to understand the psychological, social, and psychosocial implications of tailoring AI personalities to their human users and we encourage the development of that line work in future research.
-Taken together, the effects shown here underscore the critical role of personality alignment in improving human-AI collaboration.
-By demonstrating that specific pairings can improve performance while others hinder it, we highlight the nuanced dynamics at play in human-AI teams.
-The observed trade-offs between quality and productivity further illustrate the complexity of these interactions.
+Our study provides the first large scale experimental evidence that personality pairing improves human-AI collaboration and team performance and highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success.
+Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance Krakowski et al.
+(2025).
+To our knowledge, ours is the first study to measure the performance effects of AI personalization in a context in which AI personalities are directly and randomly manipulated.
+As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of performance in real agentic workflows and unlock more effective and satisfying collaborations.
+Our results extend the person-team fit literature to human-AI teams (Barrick et al., 1998, Bell, 2007, Kristof-Brown et al., 2005).
+The finding that pairings between different traits outperform same-trait pairings is consistent with complementary fit, in which team members fill each other’s gaps, rather than supplementary fit based on similarity (Kristof-Brown et al., 2005).
+The strongest effects were negative: extraverted humans produced lower quality ads when paired with conscientious AI, and conscientious humans produced lower quality ads when paired with agreeable AI.
+These results suggest that certain trait combinations create friction that impairs creative output, rather than a universal benefit of personality matching.
+The findings point to actionable pairing principles.
+For ad quality, extraverted humans produced lower-quality ads when paired with conscientious AI (the strongest effect in the study), conscientious humans produced lower-quality ads when paired with agreeable AI, and neurotic humans produced lower-quality ads when paired with conscientious AI.
+For field click-through rates, neurotic humans achieved higher CTRs when paired with neurotic AI, an effect that persisted even after controlling for ad quality.
+Moreover, the neurotic-neurotic field result aligns with how the team actually communicated.
+Among the 25 personality pairings, this was the only one that raised the pooled rate of directive language significantly ($p=0.002$, driven by both speakers), while no pairing affected message length (SI Appendix, Section LABEL:si:interaction-content).
+Our prior work (Ju and Aral, 2025) reports that directive, task-oriented communication is positively associated with click ratings in human-AI teams.
+The pairing that creates the highest increase in click-through rates is the pairing that produces the communication pattern independently shown to predict higher click through rates.
+Although we treat this convergence of results as descriptive rather than as evidence of a mediating effect, it suggests that the field effect operates through the kind of communication exchange two neurotic partners generate together and gives us greater confidence in the substantive accuracy of our results.
+While ours is the largest randomized controlled trial of AI personalization to date, several limitations circumscribe our conclusions.
+First, the specific pairing effects we identified in a marketing context may not extend to other collaborative tasks such as coding or analytics.
+More work is needed to understand AI personalization across tasks, especially given the jagged frontier of AI capabilities, which could simultaneously drive and interact with personality effects.
+Second, we randomized AI personality across the full design rather than preregistering a narrow set of targeted pairings.
+A focused preregistered replication of a small set of pairings would be an excellent next step, and in principle a stronger test of any particular pairing.
+At the same time, such a study requires knowing in advance which pairings are worth targeting.
+We therefore view our work as establishing that personality pairing can matter for human-AI collaboration and field performance, and as a foundation for future work on targeted replication and personalization optimization.
+Third, our results were obtained with a specific model snapshot (gpt-4o-2024-08-06).
+Different architectures and alignment procedures produce different behavioral responses to personality prompts, so the specific coefficients are likely model-dependent.
+The general finding that personality pairing matters should still generalize to any model capable of prompt-induced behavioral variation.
+Steering techniques themselves generalize across both models and methods, ranging from prompting to representation-level interventions (Zou et al., 2023).
+Diversifying AI personas through personality prompts has been shown to eliminate the homogenization of creative outputs observed when all users interact with identically prompted AI (Wan and Kalman, 2026).
+These prior results suggest that personality-based personalization is a robust design lever rather than an artifact of any single model or technique.
+Fourth, the Big Five personality traits may overlook other traits more relevant to human-AI dynamics.
+Our exploratory analyses point to further heterogeneity in demographic moderators like country of origin and experience with AI (SI Appendix, Section LABEL:si:demographics, and linguistic patterns in collaboration offer another promising direction for future work on dimensions relevant to AI personalization (Wan et al., 2025).
+Fifth, our scope limits what we can say about longer time horizons, larger teams, and psychosocial effects.
+A single 40-minute session cannot address how pairings play out over repeated interactions, including whether prolonged use produces over-reliance on AI (Fang et al., 2025).
+We focus on one-to-one pairings rather than multi-person teams or networks of collaborators, where AI could vary its personality across teammates or over time.
+We measure self-assessed collaboration dynamics with teamwork variables, but we examine no measures of well-being or adverse psychosocial outcomes.
+As the field of AI personalization develops, it will be important to understand the psychological and social implications of tailoring AI personalities to their human users, and we encourage future research in this area.
+While acknowledging these limitations, we believe this large-scale experimental evaluation of personality pairing in human-AI collaboration is an important first step toward robust AI personalization.
+The effects shown here underscore the critical role of personality pairing in improving human-AI collaboration.
+Specific pairings improved ad quality, teamwork, and real-world ad performance, while other pairings hindered these outcomes.
 These insights pave the way for designing AI agents that are not only technically proficient but also socially and behaviorally adaptive, ensuring they complement their human counterparts effectively.
-Ultimately, our work demonstrates that personality-aligned AI has the potential to transform collaborative workflows and performance in human-AI teams.
+Ultimately, our work demonstrates that strategically pairing AI personalities with human traits has the potential to transform collaborative workflows and performance in human-AI teams.
 Methods
 This research was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects (see osf.io/jfzha and osf.io/95dhu).
+Informed consent was obtained from all participants prior to their participation in the study.
 The Pairit platform
 The study was conducted on the Pairit platform (Figure 5).
-On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E, and write copy for the display ads.
+On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E 3, and write copy for the display ads.
 In the human-AI teams, the AI has full context of all elements on the platform, including screenshots of the ad images as they evolve over time during the collaboration.
 In all conditions, only the participants can submit ads.
-Additional details on the platform are in SI Section A-B.
+Additional details on the platform can be found in the SI.
 Figure 5: The Pairit platform.
@@
 The participant chats with an AI agent which has full context of the user interface and can modify all elements on the task panel.
+Some text in the screenshot is intentionally blurred to anonymize identifying information about the partner think tank, and a larger version of the same screenshot is available in SI Appendix, Section LABEL:si:pairit.
 To enable the AI agents to act as effective collaborators on the Pairit platform, we provided them with full context of the task and collaboration by prompting each LLM call with all information on the user interface.
 The prompts included the same text of the task given to participants, previous submissions, personality prompts, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, general instructions, and a screenshot of the image after each change by the participant.
-We used OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06).
-Additional details in the AI agent are in SI Section  D.
+We used OpenAI’s multimodal gpt-4o model (specifically the gpt-4o-2024-08-06 snapshot) with a temperature of 0.2 and default nucleus sampling parameters; no top-$p$, max token, or seed constraints were imposed, so sampling was stochastic.
+Image generation used dall-e-3 at $1024\times 1024$ resolution.
+Additional details on the AI agent are in SI Appendix, Section LABEL:si:methods.
 Using Pairit, we designed the AI agent to be able to take the same actions that a human participant could take, apart from submitting ads.
@@
 The agent is prompted every 10 seconds on whether to engage to take an action.
-To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (wei2023chainofthought).
+To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (Wei et al., 2022).
 Specifically, we prompt the model with questions to reflect on the state of the collaboration.
 These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message.
+Personality prompts
+We note that these prompts induce behavioral styles rather than latent psychological traits; we use “personality” throughout as shorthand for prompt-induced behavioral tendencies.
 For the randomization of AI personalities, we prompt the model with prompts that induce high or low levels of the Big Five personality traits.
-Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (jiang2023evaluating).
-See SI Section E for covariate balance of the randomization of each personality prompt.
+Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (Jiang et al., 2023).
+See SI Appendix, Section LABEL:si:randomization, for sample summary statistics, balance across the human-AI and human-human conditions, and the independent AI-trait randomization design.
+We validated that the personality prompts produced measurable behavioral differences in AI messages.
+For each trait, we extracted a pre-specified set of linguistic features from AI-generated messages and constructed composite behavioral indices using the sum of standardized feature values with theoretically expected signs (Anderson, 2008).
+Examples of core behavioral markers include: (1) Openness: emoji usage, conservative language (reversed), and exploratory suggestions; (2) Conscientiousness: first-message elaboration (length and word count); (3) Extraversion: enthusiasm markers, exclamation usage, and words per message; (4) Agreeableness: permission-asking ratio, affirming language, and we/our orientation; (5) Neuroticism: cautious language, fewer alternatives offered (reversed), and less confident language (reversed).
+All five composites distinguished their target trait condition (high vs.
+low) at $p<0.05$, with Cohen’s $d$ ranging from 0.12 to 0.17 (SI Appendix, Section LABEL:si:manipulation).
+Agreeableness had the strongest individual behavioral feature (permission-asking; $d=0.14$, $p=0.013$), and conscientiousness had the broadest manipulation (three individually significant features).
+A discriminant validity matrix confirmed that each composite preferentially differentiated its own trait condition, with the largest effect on the diagonal in four of five cases.
+LLMs can exhibit social desirability biases that attenuate personality manipulations (Salecha et al., 2024), which may partly explain the modest effect sizes of our behavioral composites.
 Procedure
@@
 We do not reveal whether or not the partner is a human or an AI until the post-task survey.
-Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (rammstedt2007personality).
-See SI Section E for summary statistics and covariate balance of the human-AI randomization and SI Section G for the survey items.
+Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (Rammstedt and John, 2007).
+See SI Appendix, Section LABEL:si:randomization, for summary statistics and covariate balance of the human-AI randomization, and SI Appendix, Section LABEL:si:survey, for the rating interface and ad mockups used in the evaluation study.
 Following the pairing and completion of the pre-task survey, participants entered a collaborative workspace where they could chat and collaboratively develop and submit display ads (Figure 1B).
@@
 After 40 minutes, participants were automatically re-directed to a post-task survey.
-Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (hoegl2001).
+Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (Hoegl and Gemuenden, 2001).
 The original instrument included 38 questions on a 7-point Likert scale, but three communication-related items were excluded as they were irrelevant to a one-time online collaborative effort.
@@
 Incentives
-To incentivize the production of high-quality ads, we informed the participants of their eligibility for additional \$100 as bonus payment, determined by the volume and quality of their submissions.
-Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
+To incentivize the production of high-quality ads, we informed the participants of their eligibility for an additional \$100 bonus payment, determined by the volume and quality of their submissions.
+Participants were explicitly instructed that “the greater the number of ads, the greater your chances, but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
 Human evaluation of ad quality
@@
 As a participant entered our survey platform, we drew one sample from the set, without replacement, to provide to the participant, with each participant receiving a unique sample.
-To estimate ad performance, we created and used mockups of display ads in the survey (see SI Section C).
+To estimate ad performance, we created and used mockups of display ads in the survey (SI Appendix, Section LABEL:si:survey).
 For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate.
 Each item was rated on a 7-point Likert scale.
-The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
+The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad." We construct the ad quality index at the individual rating level as the unweighted mean of each rater’s three 7-point evaluations of a given ad: text quality, image quality, and click likelihood.
 Field evaluation of ad performance
@@
 VTR and VTD were tracked via a unique DocSend link for each ad.
-Our ad campaigns assessed causal ad effects conditional on online algorithms and divergent targeting (braun2024adplatforms, braun2024abtesting).
-No holdout group was used, as unique DocSend links imply zero visits without ads (braun2024adplatforms).
-We sampled 2,000 ads from 11,138, selecting 1–2 ads per 1,834 human-human and human-AI teams using stratified sampling.
-Ads were grouped by team type and divided into 10 strata based on human-predicted click likelihood.
-Eight ads were excluded for potential policy violations (e.g., violence, sexual, or drug content).
-To avoid audience overlap, 400 campaigns each ran five unique ads as split tests.
-Each campaign targeted 133 random ZIP codes with population between 10,000 and 100,000, as per the 2020 American Community Survey.
+Ad platforms like X use auto-bidding algorithms that optimize delivery based on expected engagement, producing divergent targeting across ads within a campaign (Braun and Schwartz, 2024).
+Following Braun et al.
+(Braun et al., 2024), we estimate causal effects of ads conditional on the platform’s optimization algorithm.
+The relevant estimand is the effect of ad content on performance given the audience the platform selected for each ad, which is the economically relevant quantity for advertisers operating within algorithmic environments.
+Viewer composition is therefore not constant across ads within a campaign, but campaign random effects absorb systematic differences in audience pools across campaigns.
+Advertising spend is controlled to account for differential budget allocation from auto-bidding.
+ZIP-code-level randomization across campaigns prevents audience spillovers.
+No holdout group was used, as unique DocSend links imply zero visits without ads (Braun et al., 2024).
+The field experiment used a subset of the human-AI ads.
+We sampled 2,000 ads from the 7,266 ads produced in the human-AI lab condition using stratified random sampling based on predicted click likelihood implied by the human ratings, with 10 click-score strata spanning the ad quality distribution.
+These ads were then deployed in 400 geographically non-overlapping campaigns, each with a five-ad split test.
+The selected ads covered 51.7% of human-AI lab participants (651 of 1,258), and each included participant contributed a mean of 1.77 ads.
+Eight sampled ads were excluded for potential policy violations on the X platform (e.g., violence, sexual, or drug content).
+Each campaign targeted 133 random ZIP codes with populations between 10,000 and 100,000, as per the 2020 American Community Survey.
 One-way ANOVA yielded non-significant results across ZIP codes: $F(399,53199)=0.954$, $p=0.734$ for population; $F(399,53199)=0.973$, $p=0.636$ for income.
@@
 Campaign random effects controlled for temporal confounders.
-In our regression analyses, advertising spend was controlled for to account for auto-bidding.
+Multiple testing corrections
+Our experimental design yields 25 human-AI personality interaction terms per outcome (5 human traits $\times$ 5 AI traits).
+To account for multiple testing, we group related interaction tests into coherent outcome families and apply Benjamini-Hochberg (BH) corrections within each family.
+These families cover ad quality, teamwork, productivity (reported in the SI), and field performance, with a separate correction for each field outcome (CTR, CPC, VTR, VTD).
+We use the BH correction framework because our objective is to identify a set of statistically supported findings while controlling the expected false discovery rate among reported results, rather than to minimize the probability of a single false rejection across the entire family of tests.
+Given the number of related hypotheses evaluated, a correction framework akin to Bonferroni would impose substantially lower power and materially increase the risk of false negatives.
+Since our study is intended to characterize the pattern of supported effects rather than to certify that the probability of any false rejection is below a fixed threshold, BH is the most appropriate framework.
+We report both raw and corrected $p$-values.
+The pre-registration mapping table in SI Appendix, Section LABEL:si:prereg, indicates which analyses are treated as preregistered versus exploratory.
 Model specifications
-To evaluate the effect of personality trait interactions on human-AI team outcomes, we use the following regression model:
-$Y_{J}=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}+\epsilon_{i}\text{,}$ (1)
-where $Y_{J}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low), and $\epsilon_{i}$ is the error term.
-The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating (with standard errors clustered at the ad level); for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors).
-The outcomes of interest $Y_{i}$ include human-evaluated ad quality, productivity (i.e., the number of ads submitted), and teamwork quality measures.
+We use outcome-appropriate regression models for human-AI team outcomes.
+For ad quality and teamwork (7-point Likert scales), we estimate ordinal logit models to respect the ordinal nature of the outcome.
+For productivity (submission count), we estimate Poisson generalized linear models with heteroskedasticity-robust (HC1) standard errors to accommodate dispersion in the count distribution.
+OLS results are reported in the SI as a robustness check and are substantively similar.
+The general specification is:
+$g(E[Y_{i}])=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}\text{,}$ (1)
+where $g(\cdot)$ is the link function (identity for OLS, log for Poisson, logit for ordinal), $Y_{i}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low).
+The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating; for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors).
+All regressions are estimated within the human-AI condition, so the human-human condition is not part of the estimation sample.
+Each interaction coefficient $\beta_{ha}$ in Figure 3 therefore captures how the association between human trait $h$ and the outcome differs when the AI partner is assigned high rather than low levels of trait $a$, conditional on the human and AI main effects and the other regressors.
+For example, if $\beta_{ha}=0.5$ for $_{\text{H}} ×\times Conscientiousness$$_{\text{AI}}$, then moving from the minimum to the maximum normalized human conscientiousness score increases the model’s linear predictor by $0.5$ more when paired with high-conscientiousness AI than with low-conscientiousness AI.
 To evaluate the field experiment, we use the following mixed effects model:
@@
 where $Y_{ic}$ represents the outcome (e.g., CPC, CTR, VTR, VTD) for ad $i$ in campaign $c$, Image, Text, and Click are human-rated quality measures, $\text{Spend}_{ic}$ is the campaign spend, $u_{c}\sim N(0,\sigma^{2}_{u})$ is the random intercept for campaign $c$, and the rest are the same as in Equation 1.
+Data availability
+The data supporting the findings of this study will be made available in the paper’s data repository upon publication.
 Acknowledgments